OpenStudy (anonymous):

How to solve problems similar to 1+2/1-2 ?

OpenStudy (anonymous):

\[\frac{1+2}{1-2}=\frac{3}{-1}=-3\] what counts as similar?

OpenStudy (anonymous):

But the answer is 1.